# Learning Multiplex Embeddings on Text-rich Networks with One Text Encoder

### v1

@@ A.5.1 Multiplex Representation Learning
To facilitate the reproduction of our representation learning experiments, we provide the hyperparameter configuration in Table 8. Vanilla FT, MTDNN, and Metern use exactly the same set of hyperparameters for a fair comparison. The last layer [CLS] token hidden states are utilized to develop ${\bm{h}}_{v|r}$ for Vanilla FT, MTDNN, and Metern. Paper titles and item titles are used as text associated with the nodes in the two kinds of networks, respectively. (For some items, we concatenate the item title and description together since the title is too short.) The models are trained for 40 epochs on 4 Nvidia A6000 GPUs with a total batch size of 512. The total time cost is around 17 hours and 2 hours for networks in the academic domain and e-commerce domain respectively. Code is available at https://anonymous.4open.science/r/METER-submit-2C7B.
@@ A.5.3 Learn to Select Source Relations
Problem Definition. Given a query paper node (with associated text), we aim to predict what is the category of the paper. The number of paper node categories in academic networks (Geology and Mathematics) is shown in Table 9.
Table 9: Number of paper node categories in academic networks.
@@
Problem Definition. Given a query item node (with associated text), we aim to predict what is the category of the item. The number of item node categories in e-commerce networks (Clothes, Home, and Sports) is shown in Table 10.
Table 10: Number of item node categories in e-commerce networks.
@@
We study how different relation embedding initialization affects the quality of multiplex representations learned by Metern. We explore three initialization settings: zero vectors initialization, normal distribution initialization, and word embedding initialization. The results of average PREC@1 on different networks are shown in Table 11. From the results, there is no significant difference between the representation learning quality of different initialized relation embeddings.
Table 11: Performance of different relation embedding initialization on different networks.
@@
Metern w/ zero init 36.31 55.26 79.99 79.68 77.59
Metern w/ randn init 36.43 55.20 80.19 79.69 77.56
Metern w/ word init 36.36 55.20 80.12 79.83 77.71
@@
In section 4.2, we propose to let the model learn to select source relations for different downstream tasks and show the learned source relation weight for citation prediction and paper recommendation in Figure 3. In this section, we show more results on the learned source relation weight on academic network downstream tasks in Figure 8 and on e-commerce network downstream tasks in Figure 9. There are four downstream tasks in the academic domain: year prediction, citation prediction, paper classification, and paper recommendation. There are two downstream tasks in the e-commerce domain: price prediction and item classification. From Figure 8 and Figure 9, we can find that different relations can benefit different downstream tasks.
@@
Figure 8: Learnt source relation weights for tasks on academic networks (Geology and Mathematics). The x-axis is the learned weight vector id and the y-axis is the relation embedding id grouped by relation id. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively.
@@
Figure 9: Learnt source relation weights for tasks on e-commerce networks (Clothes, Home, and Sports). The x-axis is the learned weight vector id and the y-axis is the relation embedding id grouped by relation id. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively.
@@
We visualize the relation token embeddings $Z_{\mathcal{R}}$ learned by Metern with t-SNE Van der Maaten and Hinton [2008], projecting those embeddings into a 2-dimensional space. The results on the Geology network and Mathematics network are shown in Figure 10, where the embeddings belonging to the same relation are assigned the same color. From the results, we can find that embeddings belonging to the same relation are close to each other, while those belonging to different relations are discriminative. This demonstrates that Metern can learn to capture the different semantics of different relations by assigning different relations’ embeddings to different areas in the latent space.

### v2

@@ A.5.1 Multiplex Representation Learning
To facilitate the reproduction of our representation learning experiments, we provide the hyperparameter configuration in Table 8. Vanilla FT, MTDNN, and Metag use exactly the same set of hyperparameters for a fair comparison. The last layer [CLS] token hidden states are utilized to develop ${\bm{h}}_{v|r}$ for Vanilla FT, MTDNN, and Metag. Paper titles and item titles are used as text associated with the nodes in the two kinds of graphs, respectively. (For some items, we concatenate the item title and description together since the title is too short.) The models are trained for 40 epochs on 4 Nvidia A6000 GPUs with a total batch size of 512. The total time cost is around 17 hours and 2 hours for graphs in the academic domain and e-commerce domain respectively. Code is available at https://github.com/PeterGriffinJin/METAG.
@@ A.5.3 Learn to Select Source Relations
Problem Definition. Given a query paper node (with associated text), we aim to predict what is the category of the paper. The number of paper node categories in academic graphs (Geology and Mathematics) is shown in Table 9.
Table 9: Number of paper node categories in academic graphs.
@@
Problem Definition. Given a query item node (with associated text), we aim to predict what is the category of the item. The number of item node categories in e-commerce graphs (Clothes, Home, and Sports) is shown in Table 10.
Table 10: Number of item node categories in e-commerce graphs.
@@
We study how different relation embedding initialization affects the quality of multiplex representations learned by Metag. We explore three initialization settings: zero vectors initialization, normal distribution initialization, and word embedding initialization. The results of average PREC@1 on different graphs are shown in Table 11. From the results, there is no significant difference between the representation learning quality of different initialized relation embeddings.
Table 11: Performance of different relation embedding initialization on different graphs.
@@
Metag w/ zero init 36.31 55.26 79.99 79.68 77.59
Metag w/ randn init 36.43 55.20 80.19 79.69 77.56
Metag w/ word init 36.36 55.20 80.12 79.83 77.71
@@
In section 4.2, we propose to let the model learn to select source relations for different downstream tasks and show the learned source relation weight for citation prediction and paper recommendation in Figure 3. In this section, we show more results on the learned source relation weight on academic graph downstream tasks in Figure 8 and on e-commerce graph downstream tasks in Figure 9. There are four downstream tasks in the academic domain: year prediction, citation prediction, paper classification, and paper recommendation. There are two downstream tasks in the e-commerce domain: price prediction and item classification. From Figure 8 and Figure 9, we can find that different relations can benefit different downstream tasks.
@@
Figure 8: Learnt source relation weights for tasks on academic graphs (Geology and Mathematics). The x-axis is the learned weight vector id and the y-axis is the relation embedding id grouped by relation id. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively.
@@
Figure 9: Learnt source relation weights for tasks on e-commerce graphs (Clothes, Home, and Sports). The x-axis is the learned weight vector id and the y-axis is the relation embedding id grouped by relation id. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively.
@@
We visualize the relation token embeddings $Z_{\mathcal{R}}$ learned by Metag with t-SNE (Van der Maaten & Hinton, 2008), projecting those embeddings into a 2-dimensional space. The results on the Geology graph and Mathematics graph are shown in Figure 10, where the embeddings belonging to the same relation are assigned the same color. From the results, we can find that embeddings belonging to the same relation are close to each other, while those belonging to different relations are discriminative. This demonstrates that Metag can learn to capture the different semantics of different relations by assigning different relations’ embeddings to different areas in the latent space.
@@
A.9 More Comparison with Multiplex GNNs
We compare our methods with multiplex GNNs equipped with pretrained text embeddings (MPNet-v2 embedding). The results are shown in Figure 12 and Figure 13.
Table 12: Comparison between multiplex GNNs and Metag on Mathematics graph.
Mathematics
Model cb sa sv cr ccb Avg.
DMGI (BOW) 46.55 42.62 6.11 27.80 38.87 28.85
DMGI (MPNet) 54.13 53.06 7.40 31.39 43.98 37.99
HDMI (BOW) 52.65 52.71 5.54 31.80 42.54 37.05
HDMI (MPNet) 57.34 54.45 6.59 33.45 44.24 39.21
Ours 79.40 72.51 14.03 47.81 62.24 55.20
Table 13: Comparison between multiplex GNNs and Metag on Sports graph.
Sports
Model cop cov bt cob Avg.
DMGI (BOW) 41.37 46.27 41.24 31.92 40.20
DMGI (MPNet) 43.37 63.69 58.16 51.72 54.24
HDMI (BOW) 45.43 61.22 55.56 52.66 53.72
HDMI (MPNet) 43.12 62.65 57.88 51.75 53.85
Ours 67.92 79.85 81.52 81.54 77.71
From the result, we can find that the performance of multiplex GNN with pretrained text embeddings is consistently better than the performance of multiplex GNN with bag-of-word embeddings. However, our method can outperform both baselines by a large margin.
A.10 Comparison with Natural Language Description of Relation + LM Baselines
Another way to capture relation-specific knowledge is to use natural language descriptions of the relations and feed them into the language model for joint relation and text encoding. We implement this baseline by adding a natural language description of the relation before encoding and using the large LM OpenAI-ada-002 (best OpenAI embedding method) as the backbone text encoder. The results on the Mathematics graph and the Sports graph are shown in Figure 14 and Figure 15.
Table 14: Comparison between relation description baseline and Metag on Mathematics graph.
Mathematics
Model cb sa sv cr ccb Avg.
OpenAI-ada-002 (relation description) 38.87 26.40 2.43 30.72 18.74 23.43
Ours 79.40 72.51 14.03 47.81 62.24 55.20
Table 15: Comparison between relation description baseline and Metag on Sports graph.
Sports
Model cop cov bt cob Avg.
OpenAI-ada-002 (relation description) 49.56 71.75 58.36 54.63 58.58
Ours 67.92 79.85 81.52 81.54 77.71
From the result, we can find that our method outperforms this baseline by a large margin, this demonstrates that the relation embeddings learned by our method to represent relations are better than the natural language description used in the baseline.
A.11 Comparison with LinkBERT and GraphFormers
We compare our method with LinkBERT (Yasunaga et al., 2022) and GraphFormers (Yang et al., 2021) on the Mathematics graph and the Sports graph. The results are shown in Figure 16 and Figure 17.
Table 16: Comparison between LinkBERT, GraphFormers and Metag on Mathematics graph.
Mathematics
Model cb sa sv cr ccb Avg.
LinkBERT 74.18 62.52 8.84 44.53 59.76 49.96
GraphFormers 64.54 47.96 5.19 36.70 52.13 41.30
Ours 79.40 72.51 14.03 47.81 62.24 55.20
Table 17: Comparison between LinkBERT, GraphFormers and Metag on Sports graph.
Sports
Model cop cov bt cob Avg.
LinkBERT 67.65 76.34 80.76 77.73 75.62
GraphFormers 66.92 78.10 79.14 78.29 75.61
Ours 67.92 79.85 81.52 81.54 77.71
From the result, we can find that our method outperforms LinkBERT and GraphFormers consistently. The main reason is that these two baseline methods only generate one embedding for each text unit, expecting that all types of relations between texts can be captured by these single-view embeddings, which do not always hold in real-world scenarios. GraphFormers adopts the same graph propagation and aggregation module for edges of different semantics, which results in degraded performance.